OpenStudy (princesssleelee):

Three functions are given below: f(x), g(x), and h(x). Explain how to find the axis of symmetry for each function, and rank the functions based on their axis of symmetry (from smallest to largest).

OpenStudy (princesssleelee):

f(x) = –4(x − 8)2 + 3 g(x) = 3x2 + 12x + 15

OpenStudy (igigighjkl):

I found something.

OpenStudy (igigighjkl):

one second please.

OpenStudy (igigighjkl):

Sorry can't find it.

OpenStudy (igigighjkl):

@pooja195 can you help.

OpenStudy (igigighjkl):

Sorry.

OpenStudy (igigighjkl):

I found something.

OpenStudy (igigighjkl):

you need to convert f(x) into a quadratic equation, so f(x) = 3(x + 4)^2 + 1 f(x)=3(x^2+8x+16)+1 f(x)=3x^2+24x+49

OpenStudy (igigighjkl):

Does this help at all?

OpenStudy (igigighjkl):

@princesssleelee are you their?